Ray is now 18 years old. His mother, Kay is now 46 years old.
a)write an expression for each person's age "x" years from now.
b) Write an equation stating that Kay is twice as old as Ray. Then solve it to find out when Kay is twice as old.
c) When is Kay three times as old as Ray?
d) write an equation stating that Ray is exactly as old as Kay. Solve it. When is this true?
e) write an equation stating that Ray's age is 28 years less than Kay's age. When will this be true?

Answer by Finavon(81)

a)
Ray 18+x
Kay 46+x
b)
(18+x)*2=46+x
36+2x=46+x
x=10 years ahead
Ray 28; Kay 56
c)
(18+x)*3=46+x
54+3x=46+x
2x=-8
x=-4
4 years ago
Ray 14; Kay 42
d)
18+x = 46+x
Never true
e)
(18+x)+28=46+x
Always true